Summary: Moderate winds and partly cloudy. Hasn't rained yet which probably means it's about to. Set a course record for sites visited in a single trip by going to all but 14 to raise DSM and battery boxes a few inches for better water clearance. Placed new soil probe at S1. Took the TRH from S1 to replace the fan pcb. S10's TRH isn't reading fan current but is still working.
Actions past 24 hours:
To dos:
Station Status:
S3: Pi DSM is working.
S15: Also working.
Sensor Status:
T/RH: Lots of issues: 42m now Ifan=0, don't know if this is real;
P: ok; Pirga still lower by about 1mb
2D Gill: ok
csat u,v: S3 looked bad 2 days ago when initially restarted through Pi. Now not sure.
csat ldiag: ok
csat w, tc: ok
EC150: ok
motes: S15 was up at least a bit yesterday, but not sure when we left.
Wetness: ok
radiation: ok
Tsoil: ok
Gsoil: ok
Qsoil: S9 and S15 ok; S1 is dead (a new EC-5 sensor may be coming from Boulder)
Cvsoil: still need to have John/Steve S come up with a fix (in Boulder)
Rainr: ok
Vbatt: ok